Use 2 methods (uniform motion equations and the principle of conservation of energy) to check the accuracy of the statement at the beginning of the passage that a diver jumping from the high platform will hit the water at 30mph.


1
Expert's answer
2021-03-02T18:04:34-0500

1)


"mgh=0.5mv^2\\gh=0.5v^2\\9.8h=0.5(30\cdot0.44704)^2\\h=9.2\ m"

2)


"v^2-0^2=2gh\\"

"2(9.8)=(30\cdot0.44704)^2\\h=9.2\ m"